Theory and Applications of Categories, Vol. 32, No. 24, 2017, pp. 803–822.

METRIC SPACES AND SDG

ANDERS KOCK

Abstract. We present an axiomatic theory, based on the notions of metric space and space with a (first order) neighbour relation. The axiomatics implies a synthetic proof of Huygens’ principle of wave fronts, as envelopes of a family of spheres. A model of the axiomatics is presented in terms of synthetic differential geometry (SDG)^{1}.

Introduction

The first use of the term “Synthetic differential geometry” seems to be in Busemann’s [3],

[4] (1969 and 1970), so it is prior to the use of the term in, say, [6], where the reasoning goes in a different direction. In Busemann’s work, the basic structure is that of metric space. This notion has not been much considered in the context of SDG^{2} (except in its infinitesimal form: Riemannian metric). One reason for this is: to provide the number line R with the standard metric, one needs the absolute value function R → R, given by x ↦ |x|. This map, however, is not smooth at 0; and in SDG, only smooth maps can be considered.

However, the geometric reasoning of Busemann has a genuine synthetic character. It does admit co-existence with SDG, which I hope that the present note will illustrate.

In particular, we discuss two basic notions. The first notion is essentially the notion geodesic, or ray generated by two points; here, we follow Busemann in describing this in terms of the metric (even without assuming the triangle inequality for the distance dist, so the structure we need is only what we shall call a pre-metric). The second notion is essentially the notion of wave front in the sense of Huygens, for which, in our version, the neighbour relation, as provided by SDG, is essential. Thus the main Theorem 7.3 gives a version of the Huygens principle in synthetic terms, using the neighbour relation, and the derived notion of when two subspaces touch each other. From the touching relation, one derives the geometric notion of envelope, in terms of which the Huygens principle is formulated.

The theory developed here proceeds axiomatically, with the basic geometric axiom dealing with the touching of spheres (external and internal touching), Axioms 2.1 and 2.2, respectively. Also, two more technical sphere-touching axioms are needed, Axioms 1.7 and 1.8.

Thanks to the anonymous referee for some pertinent remarks and suggestions.

Received by the editors 2017-02-25 and, in final form, 2017-02-18.

Transmitted by Tom Leinster. Published on 2017-07-28. This version 2017-07-31.

2010 Mathematics Subject Classification: 51K10, 18B40, 18D99, 51Fxx.

Key words and phrases: metric spaces, synthetic differential geometry.

\begin{matrix}c \\ ◯ \end{matrix} Anders Kock, 2017. Permission to copy for private use granted.

1The previously published version of this article rendered pictures incorrectly on pp 805, 808 and 809.

2We shall use the acronym “SDG” when referring to the school originating with Lawvere’s “Categorical Dynamics” talk (1967) and the KL Axiom (“Kock-Lawvere”), see [6], [14], [11], [8], …

1. Metric spaces and the neighbour relation

To have a metric on a set M, one needs a number line R to receive the values of the metric, usually the ring of real numbers.

Recall that it is essential for SDG that the number line R, with its ring structure, has a rich supply of nilpotent elements, in particular, elements ϵ with ϵ^{2} = 0. Such nilpotent elements lead to the geometric notion of when two points in a space are (first order) neighbours, written x ∼ y. When the space itself is the number line R, then x ∼ y will mean (x − y)^{2} = 0.

How do such nilpotent elements coexist with the metric? We hope to demonstrate not only that they do coexist, but they enhance Busemann’s metric-based differential geometric notions, by allowing a notion for when two subspaces of M touch each other (tangency), leading to e.g. the envelopes and wave fronts, occurring already in Huygens’ work.

We shall study this axiomatically, with intended application only for the special case where the metric space M is just a Euclidean space, built on basis of the given number line R. In particular we study “lines” (or rays or geodesics) in M. Lines occur as a derived concept only. We are not using the full range of algebraic properties of R, but only the addition and order properties of the positive part R_{>0}. A model for the axiomatics are presented in Section 8; it depends on having a model for the axiomatics of SDG. Therefore, it may be that no models for the present theory exist in the category of (boolean) sets. We are really talking about interpretations and models of the theory in some topos E or other suitable category; nevertheless, we shall talk about the objects in E, as if they were just sets. This is the common practice in SDG.

1.1. A basic picture. “The shortest path between two points is the straight line.”

This may be seen as a way of describing the concept of “straight line” in terms of the more primitive concept of distance (which may be measured by time: “how long does it take to go from the one point to the other” – like “optical distance” in geometrical optics, cf. e.g. [1]).

Consider an obtuse triangle, with height ϵ at the obtuse angle (cf. figure below). In coordinates, with b as origin (0, 0), the vertices are a = (−r, 0), b′ = (0, ϵ), and c = (s, 0). The height divides the triangle in two right triangles: the one triangle has catheti of lengths r and ϵ, and the other one has catheti of lengths ϵ and s. If ϵ^{2} = 0, the length of the hypotenuse of the first triangle is then, by Pythagoras, \sqrt{r^{2} + ϵ^{2}} = \sqrt{r^{2}} = r, and the length of the hypotenuse of the other triangle is similarly s.

So the path from a to c via b′ has length r + s, just as the straight line from a to c. What distinguishes, then, in terms of length, the straight line from the path via b′?

Both have length r + s, the minimal possible length.

a  b r  b′ ϵ  c s

(1)

The answer is not in terms of extremals, but in terms of stationary or critical values ; the length of the path via b is stationary for “infinitesimal variations” (e.g. via b′), unlike the path via b′, in a sense that we shall make precise by the notion of focus: b is the focus of the set of points b′ ∼ b with distance r to a and distance s to c.

We intend here to give an axiomatic theory, involving a set M equipped with a metric dist (in a certain restricted sense we shall make precise below), and a reflexive symmetric “neighbour” relation ∼. The metric is assumed to take values in an unspecified number line R with a total strict order relation >; we will only use a few properties of the number line R, namely the additive and order-properties of R_{>0}.

The axiomatics which we present has models, built on basis of (models of) SDG; we relegate the discussion of this until the end of the paper (Section 8), to stress the fact that the theory we develop is in principle prior to any coordinatization of the geometric material.

1.2. Metric spaces. A metric space is a set M equipped with an apartness relation^{3} #, assumed symmetric, and a symmetric function dist : M ×_{#} M → R_{>0}, i.e. dist(a, b) = dist(b, a) for all a, b in M with a and b apart (here, M ×_{#} M denotes the set of (a, b) ∈ M ×M with a#b). Since dist(a, b) will appear in quite a few formulae, we use Busemann’s short notation:

dist(a, b) is denoted ab.

The triangle inequality, ac ≤ ab + bc, will play no role in the present note, except when it happens to be an equality, ac = ab + bc (which is a property that a triple of points a, b, c may or may not have). In fact, we will not be using the relation ≤ in the present note. Thus, the theory below works for what one may call pre-metric spaces, meaning that no triangle inequality is used, and it thus may possibly have applications in much more general situations. We follow Busemann in writing (abc) for the statement that the triangle equality holds for three points a, b, c (mutually apart); thus

(abc) means ab + bc = ac.

Note that (abc) implies ab < ac and bc < ac. In Busemann’s theory, (abc) is expressed verbally: “the points a, b, c are collinear (with b in between a and c)”. But (abc) will be weaker than collinearity, in our context: for, (ab′c) holds in the basic picture (1) above, but a, b′ c are not collinear. We shall below give a stronger notion [abc] of collinearity.

3one may take ”x#y” to mean ”x ≠ y, in which case it is not an added structure; however, our reasoning will not involve any negated assertions.

The sphere S(a, r) with center a and radius r > 0 is defined by

S(a, r) := {b ∈ M | ab = r}.

1.3. The neighbour relation ∼. Some uses of the (first order) neighbour relation ∼ were described in [6] §I.7, and the neighbour relation is the basic notion in [8]. Knowledge of these or related SDG texts is not needed in the following, except for where we, in Section 8, construct a model of the present axiomatics.

Objects M equipped with a reflexive symmetric relation ∼, we call spaces, for the present note. Any subset A of M inherits a space structure from M, by restriction, and with this ∼-structure, it is called a subspace of M. The spaces we consider in the present note are such subspaces of a fixed M.

The motivating examples (discussed in Section 8) are the n-dimensional coordinate vector spaces R^{n} over R, where \underline{x} ∼ \underline{y} means (y_{i} − x_{i}) · (y_{j} − x_{j}) = 0 for all i, j = 1, …, n (where \underline{x} = (x_{1}, …, x_{n}) and similarly for \underline{y}). We refer the reader to the SDG literature (notably [8]) for an exploitation of the neighbour notion for more general manifolds. A main aspect is that any (smooth) map preserves ∼. Note that in particular x ∼ 0 in R iff x^{2} = 0. It is natural in the axiomatic treatment to assume that all maps constructed preserve ∼; but to avoid vagueness, we shall rather make explicit the cases where we use this principle, namely in the proof of Lemma 3.1 (see also the last lines of Section 2), and in Proposition 7.2.

If M is furthermore equipped with a metric (or even just a pre-metric), as described in the previous Subsection, there is a compatibility requirement, namely x#y and y ∼ y′ implies x#y′; and there is an incompatibility requirement: x#y and x ∼ y are incompatible, i.e. ¬((x#y) ∧ (x ∼ y)).

It is useful to make explicit the way the (pre-)metric and the neighbour relation “interact”, in the case where M is the number line R itself, where we take the distance xy to mean |y − x| (for x#y) and take x ∼ y to mean (y − x)^{2} = 0. Note that the numericalvalue function used here is smooth on the set points x#0. For R^{n} and other spaces as a model, in the context of SDG, see Section 8 below.

Neighbours of 0 in R are in SDG called first order infinitesimals. They have no influence on the order of R; it is a standard calculation that

1.4. Proposition. If x < y, and ϵ ∼ 0, then x + ϵ < y.

Note that if we define x ≤ y to mean that “x is not > y”, then, for ϵ ∼ 0,we have that ϵ is not > 0, and similarly ϵ is not < 0. So ϵ ≤ 0, and also ϵ ≥ 0. So the relation ≤ is only a preorder, not a partial order (unless ϵ^{2} = 0 implies ϵ = 0), and so ≤ cannot in general be used to determine elements in R uniquely.

An example of the relation ∼ on a space is equality: x ∼ y iff x = y. If this is the case, we say that ∼ is trivial or that M is discrete. So the theory we are to develop for spaces with metric have as a special case (a fragment of) Busemann’s theory. The reason for introducing the ∼ relation is that it allows one to express, in geometric terms and without explicit differential calculus, the notion of a stationary (or critical) value of a function defined on M. This notion is of course related to the notion of extremal value of a function; but extremal values only make good sense in the presence of a metric, but not for a pre-metric.

We recall some notions derived from a neighbour relation ∼ on M (see also [6] I.6 and [7]): For z ∈ M, we denote by 𝔐(z) the set of z′ ∈ M with z′ ∼ z, and we call it the (first order) “monad” around z. A function (typically “distance from a given point z”) δ : M → X, defined on M, is said to have z ∈ M as a stationary value if δ is constant on 𝔐(z).

1.5. Definition. Let A and B be subsets of M, and z ∈ A ∩ B. We say that A touches B at z, (or that A and B have at least first order contact at z) if

𝔐(z) ∩ A = 𝔐(z) ∩ B.

Equivalently: for all z′ ∼ z in M, we have z′ ∈ A iff z′ ∈ B. “Touching at z” is clearly an equivalence relation on the set of subsets of M that contain z.

Note that if A touches B in z, we have 𝔐(z) ∩ A ⊆ A ∩ B and 𝔐(z) ∩ B ⊆ A ∩ B.

1.6. Definition. A subset N ⊆ M will be called focused if there is a unique n ∈ N so that n′ ∼ n for all n′ ∈ N. This unique n may be called the focus of N.

Clearly, any singleton set is focused. If ∼ is trivial (or more generally, if ∼ is transitive), then singleton subsets are the only focused subsets. Note that N being focused is a property of N, and the focus of N is not an added structure.

Two subsets A and B of M may touch each other in more than one point z. We are interested in the case where they touch each other in exactly one point z, and 𝔐(z) ∩ A (= 𝔐(z) ∩ B) is focused (then necessarily with z as focus). We then say that A and B have focused touching; in this case, we call z the touching point (note the definite article), and we call 𝔐(z)∩A = 𝔐(z)∩B the touching set of A and B. (It may be strictly smaller than A ∩ B, see Remark 8.10 below.)

In the intended application in SDG, we have, for spheres in R^{n}, with the standard Euclidean metric, the following facts, which we here take as axioms:

1.7. Axiom. Assume that A and C are spheres in M, and that b ∈ A ∩ C. Then:

𝔐(b) ∩ A ⊆ 𝔐(b) ∩ C implies 𝔐(b) ∩ A = 𝔐(b) ∩ C.

This is essentially because the spheres have the same dimension. A proof of validity of this axiom in the context of the standard SDG axiomatics is given in Proposition 8.3 below.

1.8. Axiom. If two spheres A and C in M (whose centers are apart) touch each other, then the touching is focused.

The validity of this Axiom in the intended model for the axiomatics is argued in Section 8 below (Proposition 8.9).

The following gives a characterization of the focus asserted in Axiom 1.8. Let A and C be as in the Axiom.

1.9. Proposition. Assume b ∈ A ∩ C, and assume that for all b′, we have

(b′ ∼ b ∧ b′ ∈ A) ⇒ b′ ∈ C.  (2)

Then b is the touching point of A and C.

Proof. The assumption (2) gives 𝔐(b) ∩ A ⊆ 𝔐(b) ∩ C, and then Axiom 1.7 gives 𝔐(b) ∩ A = 𝔐(b) ∩ C. So A and C touch at b. Since the touching is focused, by Axiom 1.8, the term “the touching point” is justified.

2. Touching of spheres

Let M be any (pre-)metric space, with a neighbourhood relation ∼, as in 1.2 and 1.3. For two spheres in M, one has two kinds of touching, external and internal. External touching occurs when the distance between the centers equals the sum of the radii, and internal touching when the distance between the centers is the (positive) difference between the radii. In elementary Euclidean geometry, the differential-geometric concept of “touching” may, for spheres, be replaced by the more primitive concept of “having precisely one point in common”, so classical synthetic geometry circumvents bringing in differential calculus for describing the touching of two spheres. In our context, the use of differential calculus is replaced by using the notion of touching derived from the synthetic neighbour relation ∼, as described in Section 1.3; it is applicable to any two subspaces of M. The classical criteria for touching of spheres in terms of the distance between their centers then look the same as the classical ones, except that the meaning of the word “touching” is now the one defined using ∼. These criteria we take as axioms:

2.1. Axiom. [External touching] Let A = S(a, r) and let C = S(c, s) with ac > r. Then the following conditions are equivalent:

1) A and C touch each other

2) ac = r + s.

The touching point of A and C implied by 1) and Axiom 1.8 is denoted b in the following picture. We shall use the notation a◁_{s}c for b; r need not be mentioned explicitly, it is ac − s.

A  b C  “b = a ◁_{s} c”  (3)

Using Proposition 1.9, this b may be characterized by

for all b′ ∼ b : ab′ = ab ⇒ b′c = bc  (4)

and also by

for all b′ ∼ b : b′c = bc ⇒ ab′ = ab.  (5)

Since b ∈ A ∩ C, we have ab = r and bc = s, and since we also have ab + bc = ac (by condition 2 in the Axiom), we have the triangle equality ab + bc = ac; recall the notation (abc) for this equality.

2.2. Axiom. [Internal touching] Let A = S(a, r + s) and let B = S(b, s) with ab < r + s. Then the following conditions are equivalent:

1) A and B touch each other

2) ab = r.

(Note that the sphere A here is not the same as A in the previous Axiom; it is bigger.) The touching point of A, B implied by 1) and Axiom 1.8 is denoted c in the following

picture. We shall use the notation a ▷_{s} b for c; again r need not be mentioned explicitly.

A  B c  “c = a ▷_{s} b”  (6)

Using Proposition 1.9, this c may be characterized by

for all c′ ∼ c : ac′ = ac ⇒ bc′ = bc  (7)

and also by

for all c′ ∼ c : bc′ = bc ⇒ ac′ = ac.  (8)

Again, we have the triangle equality ab + bc = ac.

2.3. Interpolation and extrapolation. We shall describe how the Axioms for external and internal touching of spheres give rise to an interpolation process and to an extrapolation process, respectively.

More precisely, given two points a and c (with a#c) and given a number s with 0 < s < ac. Consider the two spheres

S(a, ac − s) and S(c, s).

The sum of the two radii is ac, so the Axiom for external touching states that the touching of S(a, ac−s), S(c, s) is focused. Denote the touching point by a◁_{s}c. This is the b depicted in (3), with r = ab, s = bc. Note that ab + bc = ac, or in Busemann’s notation (abc).

Also, let there be given two points a and b, and an arbitrary number s > 0. Consider the two spheres

S(a, ab + s) and S(b, s).

The difference of the two radii is ab, so the Axiom for internal touching states that the touching of S(a, ab + s), S(b, s) is focused. Denote the touching point by a ▷_{s} b. This is the c depicted in (6) with r = ab, s = bc. Note that we also here have (abc).

The notation a ◁_{s} c suggests that a ◁_{s} c is the point obtained by moving s units from c in the direction given by the “vector” from c to a; it is an interpolation, since b is in between a and c, by (abc). Likewise a ▷_{s} b is the point obtained by moving s units away from b in the direction given by the “vector” from a to b; it is an extrapolation.

The possibility of extrapolation is a basic axiom in Busemann’s synthetic geometry, Axiom D in [2], Section II. Geometrically, this axiom says that any line segment from a point a to another point b may be extrapolated (prolonged) beyond b by the amount of s units say, for certain s ∈ R_{>0}. The theory we present makes a more rude statement about extrapolation, namely that extrapolation for any positive amount s is possible, and this implies that the spaces we consider are unbounded. (Busemann was also interested in bounded models for his axiomatics, namely e.g. elliptic spaces.)

We note that we have constructed a map b ↦ a ▷_{s} b from S(a, r) to S(a, r + s), (one should think of it as radial projection for two concentric spheres); as any map that can be constructed, we assume that it preserves ∼: if b_{1} ∼ b_{2} in S(a, r), then a ▷_{s} b_{1} ∼ a ▷_{s} b_{2} in S(a, r + s).

3. Collinearity

The triangle equality ab + bc = ac, or (abc), for three points a, b, c, is central in the synthetic differential geometry of Busemann, for defining geodesics, and in particular lines. It expresses classically a collinearity property of a, b, c. In the present version of SDG, based on the neighbour relation, (abc) is weaker than collinearity; referring to the “basic picture” (1), we do have (ab′c), but a, b′, c will not be collinear in the stronger sense to be presented; but (again referring to the basic picture), a, b, c will.

The equivalent conditions of the following Lemma will serve as definition (Definition 3.2 below) of when three points a, b, c satisfying the triangle equality (abc) deserve the name of being collinear in our stronger sense:

3.1. Lemma. Let there be given three points a, b, c, mutually apart, satisfying (abc), i.e. satisfying the triangle equality ab + bc = ac. Then the following six assertions are equivalent:

a1 : for all a′ ∼ a : a′b = ab ⇒ a′c = ac
a2 : for all a′ ∼ a : a′c = ac ⇒ a′b = ab
b1 : for all b′ ∼ b : ab′ = ab ⇒ b′c = bc
b2 : for all b′ ∼ b : b′c = bc ⇒ ab′ = ab
c1 : for all c′ ∼ c : ac′ = ac ⇒ bc′ = bc
c2 : for all c′ ∼ c : bc′ = bc ⇒ ac′ = ac

Proof. We note that b1 and b2 are equivalent: they both express that b is the touching point of S(a, r), S(c, s) (where r = ab and s = bc), as we observed in (4) and (5), i.e. they express b = a ◁_{s} c. Similarly c1 and c2 are equivalent: they both express that c is the touching point c in of S(a, r + s) and S(b, s), i.e. c = a ▷_{s} b. Finally, a1 and a2 are equivalent, using a change of notation and the equivalence of c1 and c2.

We use b1 to prove c1. Given c′ ∼ c with ac′ = ac (= r + s). Let b′ be a ◁_{s} c′, so ab′ = r. And b′ ∼ b, since a◁_{s} preserves ∼; furthermore, b′ is characterized by

b’1: for all b′′ ∼ b′ we have ab′′ = r implies b′′c′ = s.

Then since b ∼ b′ and ab = r, we use b’1 with b′′ = b to conclude bc′ = s.

Similarly, we use c1 to prove b1: Given b′ ∼ b with ab′ = r. Let c′ be a ▷_{s} b′, so ac′ = r + s. And c′ ∼ c since a▷_{s} preserves ∼; furthermore c′ is characterized by

c’1: for all c′′ ∼ c′ we have ac′′ = r + s implies b′c′′ = s.

Then since c ∼ c′ and ac = r + s, we use c’1 with c′′ = c to conclude b′c = s. The remaining implications are proved by the same method.

3.2. Definition. Let there be given three points a, b, and c, mutually apart. Then we say that a, b, c are collinear (with b in between a and c), and we write [abc], if (abc) holds, and one of the six equivalent conditions of Lemma 3.1 holds.

Note that for bc = s, the assertion [abc] is equivalent to b = a◁_{s}c, and also to c = a▷_{s}b. Thus

3.3. Proposition. Assume that a, b, and s are given. Then the point c = a ▷_{s} b is characterized by bc = s and the collinearity condition [abc]. Also, given a, c, and s with s < ac; then the point b = a◁_{s} c is characterized by the same two conditions. In particular, b = a ◁_{s} (a ▷_{s} b), and, for s < ac, c = a ▷_{s} (a ◁_{s} c).

Geometrically, the last assertion in the Proposition just describes the bijection between the two concentric spheres S(a, r) and S(a, r + s) which one obtains by radial projection from their common center a.

Sometimes, we shall write [abc]_{1} to mean that [abc] holds by virtue of a1 or a2, and [abc]_{2} if it holds by virtue of b1 or b2, and [abc]_{3} if it holds by virtue of c1 or c2, respectively. We clearly have

[abc]_{1} iff S(b, ab) touches S(c, ac) in a,  (9)
[abc]_{2} iff S(a, ab) touches S(c, bc) in b,  (10)
[abc]_{3} iff S(a, ac) touches S(b, bc) in c.  (11)

Collinearity is “associative”, in the following sense. Given a list of four points a, b, c, d, mutually apart. Consider the following four collinearity assertions:

[abc], [abd], [acd], [bcd].

3.4. Proposition. If two of these collinearity assertions hold, then they all four do.

Proof. The proofs of the various cases are similar, so we give just one of them: we prove that [abc] and [acd] imply [bcd]. The point c occurs in all three of these assertions, and we concentrate on that point: we use [abc]_{3} and [acd]_{2} to prove [bcd]_{2}. So assume that c′ ∼ c with bc′ = bc. By [abc]_{3}, we therefore have ac′ = ac. By [acd]_{2} we therefore have the desired c′d = cd. This proves [bcd]_{2}.

Note that since ab = ba etc., the assertion [cba]_{1} is the same as [abc]_{3}. From this we conclude that collinearity is symmetric: [abc] iff [cba]. We say that three points a, b, and c are aligned if some permutation of them are collinear (so for the term “alignment”, we ignore which point is in the middle).

3.5. Proposition. Assume that [a′ab]. Then a′ ▷_{s} b = a ▷_{s} b.

Proof. By construction, a′▷_{s}b is aligned with a′ and b, and a′, a, b are aligned by assumption. So we have two of the four possible alignment assertions for a′, a, b, and a′ ▷_{s} b. From associativity of collinearity (Proposition 3.4) we conclude that a, b, and a′ ▷_{s} b are aligned (with b in the middle); and a′ ▷_{s} b has distance s to b. These two properties characterize a ▷_{s} b by Proposition 3.3.

The characterization of a ▷_{s} b (respectively of a ◁_{s} c) implies

3.6. Proposition. If two spheres touch another, then their centers are aligned with their touching point.

3.7. Stiffness. A major aim in Busemann’s work, and also in the present note, is to construct a notion of line in terms of distance. Basic here is the notion of when three a, b, c points are collinear; classically, this is the statement (abc), i.e. ab + bc = ac; this is, in practical terms, to define lines in terms of taut strings^{4}. This gives you something which is only rigid “longitudinally”, but not “transversally”, whereas our stronger notion, defined using [abc], further involves transversal rigidity: if [abc] holds, then the infinitesimal transversal variation given by replacing b by b′, as in the basic picture (1), still satisfies (ab′c), (whereas [ab′c] fails).

In practical terms, (abc) refers to lines given by a taut string, whereas [abc] refers to lines given by a ruler (or straightedge). The transversal rigidity, usually called its stiffness, of a ruler, is achieved by the width of the ruler; the stiffness makes the ruler better adapted than strings for drawing lines, when producing technical drawings on paper.

The stiffness of [abc] is obtained by a qualitative (infinitesimal) kind of width, given by the neighbour relation.

4. The ray given by two points

The notion of ray to be given now is closely related to what [2] calls a geodesic, except that a geodesic in M is (represented by) a map R → M, whereas a ray is a map R_{>0} → M, so is only a “half geodesic”; and, furthermore, a ray has, unlike a geodesic, a definite source or starting point.

4.1. Proposition. Let a and b in M, with ab = r, say. Then for any s, t ∈ R_{>0}, we have

a ▷_{t} (a ▷_{s} b) = a ▷_{s+t} b.

4the word “line” in geometry is derived from “line” (thread made of linen) in textiles.

Proof. Let for brevity c := a ▷_{s} b and d := a ▷_{t} c. Then [abc] and [acd], hence by Proposition 3.4, we also have [abd] and [bcd], Also, by construction, bc = s and cd = t. By [bcd] we have bd = s + t, and by [abd], d is aligned with a, b. These two properties characterize a ▷_{s+t} b.

We call the map R_{>0} → M given by s ↦ a ▷_{s} b the ray generated by a and b, and we call b its source of the ray. (Note that we cannot say a ▷_{0} b = b, since a ▷_{s} b only is defined for s > 0. However, it is easy to “patch” rays, using Proposition 3.4.)

4.2. Proposition. Any ray R_{>0} → M is an isometry i.e. is distance preserving. Furthermore, any triple of mutually apart points on a ray are aligned.

Proof. The first assertion is an immediate consequence of Proposition 4.1; the second follows from the collinearity of a, b, and a ▷_{s} b by iterated use of use of Proposition 3.4.

Thus, a ray with source b can be viewed as a “parametrization of its image by arc length, measured from b” (except that we have not attempted to define these terms here). Note that b itself is not in the image of the ray.

The following Example refers to the model of the axiomatics which one obtains from SDG, as in Section 8. It shows that the isometry property is not sufficient for being a ray:

4.3. Example. Consider in R^{2} the points a = (−1, 0), b = (0, 0), and consider the ray s ↦ a ▷_{s} b; it is, of course, the positive x-axis, i.e. the map s ↦ (s, 0). But for any ϵ with ϵ^{2} = 0, the map given by s ↦ (s, ϵ · s^{2}) has the isometry property expressed by (xyz) for any three values corresponding to s_{1} < s_{2} < s_{3}; but the map is not a ray, since we cannot conclude [xyz] unless ϵ = 0.

5. Huygens’ Theorem for spheres

Let T be a space, and let S_{t}, for t ∈ T, be a family of subspaces of a space M.

An envelope (note the indefinite article “an” ) for the family S_{t} (t ∈ T) is a space E ⊆ M such that every S_{t} touches E, and every point in E is touched by an S_{t}. (Here, we used the impredicative, or implicit, definition of the notion of envelope. See [7] for a comparison with a more explicit definition, equivalent to the “discriminant” method, which provides the (maximal) envelope, as the union of the “characteristics”).

5.1. Theorem. [Huygens] An envelope E of the S(b, s), as b ranges over S(a, r), is S(a, r + s). For b ∈ S(a, r), E touches S(b, s) in a ▷_{s} b.

Proof. For b ∈ S(a, r), S(b, s) touches S(a, r +s) in a▷_{s} b. Conversely, let c ∈ S(a, r +s); we take b := a ◁_{s} c. The point b is then in S(a, r), by construction. So S(b, s) touches S(a, r + s) in a ▷_{s} b, but since b = a ◁_{s} c, this is a ▷_{s} (a ◁_{s} c), which is c, by Proposition 3.3.

6. Contact elements

The notion of contact element is trivial if ∼ is discrete, for then contact elements are just one-point sets. A one point set does not generate a “ray orthogonal to it”, as the contact elements, which we are to consider, do.

6.1. Definition. A contact element at b ∈ M is a subset P of M which may be written in the form 𝔐(b) ∩ A, for some sphere A with b ∈ A.

Let A, b and P be as in the definition. We then say that A touches P at b. If C is a sphere touching A at b, we have

𝔐(b) ∩ C = 𝔐(b) ∩ A = P.

Note that P is a focused set, with focus b.

For any sphere A and any b ∈ A, there exists (many) spheres C touching A at b. (This follows by applying extrapolation and interpolation). We therefore may equivalently describe a contact element at b as the touching set of two spheres, touching another at b.

In Subsection 6.4, we will refine the notion into that of a transversally oriented contact element.

In the intended applications, the contact elements in M make up the total space of the projectivized cotangent bundle of M.

6.2. Definition. Let there be given a contact element P at b ∈ M, and let c#b. We say that c is orthogonal to P, written c ⊥ P, if for all b′ ∈ P, b′c = bc.

(Thus, if ∼ is trivial, then all points c#b are orthogonal to P.) We clearly have c ⊥ P iff P ⊆ S(c, s) (where s denotes bc). Equivalently, if a sphere S(c, s) touches P, we have c ⊥ P.

6.3. Proposition. Assume that [abc] holds. Let P be a contact element at a. Then b ⊥ P implies c ⊥ P (and vice versa).

Proof. Assume b ⊥ P. Let r denote ab and s denote bc, so ac = r + s. For any a′ ∼ a, a′b = r implies a′c = r + s, by the [abc]-assumption (in the manifestation a1 in Lemma 3.1). Also P ⊆ 𝔐(a). Since a′b = r for all a′ ∈ P, we therefore have a′c = r + s for all a′ ∈ P, which is the condition c ⊥ P. The other implication is similar.

Similarly, if [abc] and if P is a contact element at b, we have that a ⊥ P iff c ⊥ P. Finally, if [abc] and if P is a contact element at c, we have that a ⊥ P iff b ⊥ P.

6.4. Transversal orientation of contact elements. Let there be given a contact element P. The set of spheres touching P falls in two classes: two such spheres are in the same class if they touch another internally. To provide a contact element P with a transversal orientation means to select one of these two classes of spheres; the selected spheres we describe as those that touch P on the negative side (we also say: on the inside).

Just as a contact element at c may be presented as the touching set of any two spheres which touch each other at c, a transversally oriented contact element at c may be presented as the touching set of any two spheres touching another, from the inside, at c.

Let P be a contact element at b, and let c ⊥ P and bc = s. Then the sphere S(c, s) touches P. Let P be equipped with a transversal orientation; then we say that c is on the positive side of P if the sphere S(c, s) touches P from the outside.

6.5. The ray given by a contact element. Recall that c = a ▷_{s} b is characterized by bc = s and [abc]. This gives rise to another characterization of c = a ▷_{s} b in terms of ⊥: let P be the transversally oriented contact element 𝔐(b) ∩ S(a, r), where S(a, r) touches P on the inside. Then: if c ⊥ P with c on the positive side of P, and bc = s, then c = a ▷_{s} b. This follows from Proposition 3.3 (with s = bc).

Let there be given a transversally oriented contact element P at b, and an s > 0. We shall describe a point P ⊢ s by the following procedure: pick a sphere A = S(a, r) touching P from the inside (so r = ab), so [abc] with c = a ▷_{s} b. It follows from Proposition 3.5 that this only depends on the s and the transversally oriented contact element P, but not on any particular choice of the sphere A touching P from the inside; we put P ⊢ s := a ▷_{s} b. The notation suggests graphically the fact that this point c = P ⊢ s is characterized by being on the positive side of P and satisfying bc = s and c ⊥ P. The ray generated by P is defined by s ↦ P ⊢ s.

6.6. Inflation of spheres. Let there be given a sphere S(a, q), and a t > 0. The t-inflation (or the t-dilatation) of this sphere is by definition the sphere S(a, q + t).

6.7. Proposition. If two spheres touch each other internally, the two t-inflated spheres likewise touch each other internally. If the two first spheres are S(a, r + s) and S(b, s), respectively, with touching point c, then the touching point of the two t-inflated spheres S(a, r + s + t), S(b, s + t) is

a ▷_{s+t} b = a ▷_{t} (a ▷_{s} b) = a ▷_{t} c = b ▷_{t} c.

Proof. The proof of the touching assertion is identical to the classical proof, using that internal touching of spheres is equivalent to: difference of the radii equals distance between centers; for our notion of touching, this is Axiom 2.2. For the second assertion: By definition of the ▷-construction, the first expression here is the touching point of the inflated spheres; it equals the next expression by Proposition 4.1. It in turn equals the third expression, since c = a ▷_{s} b by construction. Finally, the fourth expression follows by Proposition 3.5 from collinearity of a, b and c.

6.8. Flow of contact elements. Let there be given a transversally oriented contact element P at b, as in Subsection 6.4. The construction of the ray s ↦ P ⊢ s given there can be enhanced to a parametrized family of transversally oriented contact elements s ↦ P ⊩ s, with P ⊢ s as focus of P ⊩ s. Pick, as in Subsection 6.4, a sphere A = S(a, r) touching P from the inside at b. The inflated sphere S(a, r + s) contains a ▷_{s} b = P ⊢ s, so we get a contact element 𝔐(P ⊢ s) ∩ S(a, r + s), and we take this as P ⊩ s. We orient it by declaring that S(a, r + s) touches it from the inside. We have to see that the construction is independent of the choice of A. We know already that the focus P ⊢ s is independent of the choice. If we had chosen another A′ to represent P, the spheres A′ and A touch each other from the inside at b, hence their s-inflated versions likewise touch each other from the inside, at P ⊢ s, by Proposition 6.7. This means that they define the same contact element at this point.

We leave to the reader to prove

6.9. Proposition. For P a transversally oriented contact element,

P ⊩ (s + t) = (P ⊩ s) ⊩ t,

and also

(P ⊩ s)^{∗} ⊩ s = P,

where (P ⊩ s)^{∗} denotes P ⊩ s but with opposite orientation.

Another description of the contact element P ⊩ s, again only seemingly dependent on the choice of the sphere A = S(a, r), is

P ⊩ s := 𝔐(a ▷_{s} b) ∩ S(a, r + s) = 𝔐(a ▷_{s} b) ∩ S(b, s).  (12)

7. Huygens’ Theorem for hypersurfaces

7.1. Definition. A hypersurface in M is a subset B ⊆ M which satisfies: for every b ∈ M, 𝔐(b) ∩ B is a contact element. To give such B a transversal orientation is to give every such contact element a transversal orientation.

For suitable s > 0, we aim at describing “the parallel surface to B at distance s (in the positive direction)”.

Consider a transversally oriented hypersurface B. For each b ∈ B, we have a transversally oriented contact element B(b) := 𝔐(b) ∩ B, and therefore we have the ray which it generates.

If a point x ∈ M has x ⊥ B(b), one says that b is a foot of x on B. A given x may have several feet on B; thus if for instance x is the center of a sphere B, then every point b ∈ B is a foot of x on B.

Now consider, for a given b ∈ B, the ray generated by the transversally oriented contact element B(b). Every point x on this ray has b as a foot on B. We assume that for sufficiently small s, b is the unique foot of B(b) ⊢ s on B. Consider a fixed such s. Denote the set of points c, obtained as B(b) ⊢ s for some b ∈ B, by C (or B ⊢ s). Thus we have a bijection B → C; b ∈ B and c ∈ C correspond under the bijection if b is the foot of c on B (equivalently, if c = B(b) ⊢ s). Then C deserves the name parallel surface to B, as will be argued.

We have to make the following assumption, which in the intended application is a weak one: if a point y has a unique foot on B, then so does any x ∼ y, and the two feet are neighbours.

7.2. Proposition. Under this assumption: if b ∈ B and c ∈ C correspond, then 𝔐(c) ∩ S(b, s) = 𝔐(c) ∩ C. In particular, C is a hypersurface, and is an envelope of the S(b, s), as b ranges over B. Also, 𝔐(c) ∩ C = B(b) ⊩ s.

Proof. Let x ∈ 𝔐(c) ∩ S(b, s). Let b′ be the foot of x on B. Since x ∼ c, b′ ∼ b. Since bx = s, and x ⊥ B(b′), we therefore have b′x = s. But, together, x ⊥ B(b′) and b′x = s characterize the point on C corresponding to b′; so x ∈ C.

Conversely, let c′ ∈ C and c′ ∼ c. Then c′ corresponds to a point b′ ∈ B with b′ ∼ b, implying that c′ = B(b′) ⊢ s, hence b′c′ = s; since c′ ⊥ B(b′) and b ∈ B(b′), we have bc′ = s; so c′ ∈ S(b, s). So 𝔐(c) ∩ S(b, s) = 𝔐(c) ∩ C.

Since 𝔐(c) ∩ S(b, s) is a contact element for every c ∈ C, it now follows that C is a hypersurface. Since 𝔐(c) ∩ C = 𝔐(c) ∩ S(b, s), the last assertion follows by (12) using c = a ▷_{s} b.

Recall that the C of this Proposition was more completely denoted B ⊢ s, and it deserves the name of “hypersurface parallel to B at distance s”. It inherits a transversal orientation from that of B.

We have therefore a generalization of Huygens’ Theorem, stated in [1] p. 250. The surfaces B ⊢ s mentioned are the “wave fronts”, or the “dilatations” of B ([13] p. 14-15). The Huygens Theorem stated in Section 5 is the special case where B = S(a, r).

7.3. Theorem. Let there be given a transversally oriented hypersurface B in M. Then for small enough s, we have another hypersurface B ⊢ s, which is an envelope of the spheres S(b, s) as b ranges over B. We have B ⊢ (s + t) = (B ⊢ s) ⊢ t, for t and s small enough.

The last assertion follows from Proposition 4.1, together with the characterization of P ⊢ s in terms of a▷_{s} (Subsection 6.5).

7.4. Remark. Let us note that if b ∼ b′ ∈ B, then the two contact elements P := 𝔐(b) ∩ B and P′ = 𝔐(b′) ∩ B are in united position: we say that two contact elements P and P′ are in united position if they are neighbours in the space of contact elements, and if b ∈ P′ and b′ ∈ P; this notion plays a central role in the work of S. Lie, [13] p. 39, or [12] p. 480.

The construction of rays given by “vectors” (pairs a and b of points) should be contrasted with the construction of the “flow” of contact elements P; this is in some sense the relationship between the Lagrangian and the Hamiltonian description of the process of propagation in geometrical optics, as in [1]. The present note began as an attempt to complete the essentially synthetic/metric account of this relationship, given in loc.cit. p. 250.

8. Models based on SDG

We consider in the present Section models for dist, #, and <, which are built from a (commutative) local ring R with a strict total order <. So the set of invertible elements in R fall in two disjoint classes R_{>0} and R_{<0}, both stable under addition, and with R_{>0} stable under multiplication and containing 1. We have x > y if (x − y) ∈ R_{>0}, and x < y if (x − y) ∈ R_{<}. For x − y invertible, one has the dichotomy: x < y or x > y. We write x#y for y − x invertible. The absolute value function is the function R_{<0} ∪ R_{>0} → R_{>0} given by x ↦ −x if x < 0 and x ↦ x if x > 0.

We require in the present Section, that positive square roots of elements in R_{>0} exist uniquely.

If \underline{x} = (x_{1}, …, x_{n}) ∈ R^{n} has at least one of the x_{i}s invertible, we say that \underline{x} is a proper vector. For a proper vector \underline{x}, we ask that ∑ x_{i}^{2} is > 0, so \sqrt{∑ x_{i}^{2}} ∈ R_{>0} exists. So for a proper vector \underline{x} we may define |\underline{x}| := \sqrt{∑ x_{i}^{2}}, equivalently, using the canonical inner product ⟨−, −⟩,

|\underline{x}|^{2} = ⟨\underline{x}, \underline{x}⟩.

We say that vectors \underline{x} and \underline{y} in R^{n} are apart (written \underline{x}#\underline{y}) if \underline{y} − \underline{x} is a proper vector; then |\underline{y} − \underline{x}| ∈ R_{>0} defines a metric dist(\underline{x}, \underline{y}), the distance between \underline{x} and \underline{y}.

We take for ∼ on R^{n} the standard one from SDG, namely

(x_{1}, …, x_{n}) ∼ (y_{1}, …, y_{n}) if (x_{i} − y_{i}) · (x_{j} − y_{j}) = 0 for all i, j = 1, … n.

For \underline{a} ∈ R^{n} and r > 0, the sphere S(\underline{a}, r) is the set of \underline{x} ∈ R^{n} with ⟨\underline{a} − \underline{x}, \underline{a} − \underline{x}⟩ = r^{2}. If 0 ∈ S(\underline{a}.r), we thus have ⟨\underline{a}, \underline{a}⟩ = r^{2}. The monad 𝔐(0) is D(n); elements \underline{d} in D(n) satisfy ⟨\underline{d}, \underline{d}⟩ = 0 (but ⟨\underline{d}, \underline{d}⟩ = 0 does not imply \underline{d} ∈ D(n) unless n = 1). Consider also the hyperplane H = \underline{a}^{⊥} ⊆ R^{n} (where \underline{a} is a proper vector). Then

D(n) ∩ S(\underline{a}, r) = D(n) ∩ H;

for if \underline{d} ∈ D(n) ∩ S(\underline{a}, r), we have ⟨\underline{a} − \underline{d}, a − \underline{d}⟩ = r^{2}, and if we calculate the left hand side here, we get

⟨\underline{a}, \underline{a}⟩ − 2⟨\underline{d}, \underline{a}⟩ + ⟨\underline{d}, \underline{d}⟩ = r^{2} − 2⟨\underline{d}, \underline{a}⟩,

and this can only be r^{2} if ⟨\underline{d}, \underline{a}⟩ = 0, so \underline{d} ∈ \underline{a}^{⊥} = H. Conversely, if \underline{d} ∈ D(n) ∩ H, the same calculation (essentially ”Pythagoras”) shows that \underline{d} ∈ S(\underline{a}, r).

Since the metric is invariant under translations, we therefore also have

8.1. Proposition. For A a sphere with center \underline{a}, and for \underline{b} ∈ A, we have 𝔐(\underline{b}) ∩ A = 𝔐(\underline{b}) ∩ H, where H denotes the hyperplane orthogonal to \underline{b} − \underline{a} through \underline{b}.

The following Proposition depends on R being a model for the KL axiomatics. We shall use coordinate free notation. In particular, for an n-dimensional vector space V, we have a subset D(V) ⊆ V, defined as the image of D(n) ⊆ R^{n} under some linear isomorphism R^{n} → V; it does not depend on the choice of such isomorphism.

8.2. Proposition. Let H and K be (affine) hyperplanes in an n-dimensional vector space V, and assume \underline{b} ∈ H ∩ K. Then 𝔐(\underline{b}) ∩ H ⊆ K implies H = K.

Proof. Again by parallel translation, we may assume that \underline{b} = \underline{0}, so the hyperplanes H and K are linear subspaces of V of dimension n − 1. So for dimension reasons, it suffices to prove H ⊆ K. Let ϕ : V → R be a surjective linear map with kernel K. To prove H ⊆ K, we should prove that ϕ annihilates H. By assumption, ϕ annihilates D(V) ∩ H. Since H is a linear retract of V, D(V) ∩ H = D(H) (see the proof of Proposition 1.2.4 in [8]). So the linear map ϕ |_{H}: H → R restricts to the zero map on D(H). By the KL axiom (see e.g. [8], I.3, the zero map is the only linear map which does so, so ϕ restricts to 0 on H.

Consider two spheres A and C, with centers are \underline{a} and \underline{c}, respectively, and with \underline{a}#\underline{c}. Let \underline{b} ∈ A ∩ C.

8.3. Proposition. If 𝔐(\underline{b}) ∩ A ⊆ C, then 𝔐(\underline{b}) ∩ A = 𝔐(\underline{b}) ∩ C.

Proof. Let H be the hyperplane associated to A, \underline{b} as in Proposition 8.1, and let K similarly be the hyperplane associated to C, \underline{b}. So

𝔐(\underline{b}) ∩ H = M(b) ∩ A ⊆ 𝔐(\underline{b}) ∩ C = 𝔐(\underline{b}) ∩ K.

Then Proposition 8.2 gives that H = K, and therefore the middle equality sign in

𝔐(\underline{b}) ∩ A = 𝔐(\underline{b}) ∩ H = 𝔐(\underline{b}) ∩ K = 𝔐(\underline{b}) ∩ C.

8.4. Proposition. For any \underline{x} ∈ R^{n}, the monad 𝔐(\underline{x}) is focused, with \underline{x} as focus.

Proof. For simplicity of notation, we prove that the monad 𝔐(\underline{0}) = D(n) is focused. Now D(n) may be described as \{\underline{d} = (d_{1}, …, d_{n}) ∈ R^{n} | d_{i} · d_{j} = 0 for all i, j \}. For D(1), one writes just D; so d ∈ D means d^{2} = 0. So assume that \underline{x} = (x_{1}, …, x_{n}) ∈ D(n) has \underline{x} ∼ \underline{d} for all \underline{d} ∈ D(n). This means that for all i, j, we have 0 = (x_{i} − d_{i}) · (x_{j} − d_{j}); but

(x_{i} − d_{i}) · (x_{j} − d_{j}) = −x_{i} · d_{j} − x_{j} · d_{i},

using x_{i} · x_{j} = 0 and d_{i} · d_{j} = 0. Take in particular \underline{d} of the form (d, 0, …, 0), with d ∈ D and take i = j = 1. Then the equation gives for all d ∈ D that 0 = −2x_{1} · d, so for all d ∈ D, we have x_{1} · d = 0. By cancelling the universally quantified d, we get x_{1} = 0, by the basic axiom for SDG, see [6] I.1. Similarly, we get x_{2} = 0 etc., so \underline{x} = \underline{0}.

More generally, one may prove that for suitable subspaces H ⊆ R^{n}, the set 𝔐(\underline{z}) ∩ H is focused (for \underline{z} ∈ H). This applies e.g. to a hyperplane H (zero set of a proper affine map R^{n} → R); for, then 𝔐(\underline{z}) ∩ H is ≅ D(n − 1). It also applies to spheres, which is our main concern.

A subset N of a monad 𝔐(\underline{x}), with \underline{x} ∈ N, need not be focused. Consider for example some ϵ with ϵ^{2} = 0, and consider the set \{ϵ · \underline{x} | \underline{x} ∈ R^{n}\}. It is a subset of 𝔐(\underline{0}) = D(n) and contains \underline{0}. But it is not focused: any pair of elements ϵ·\underline{x} and ϵ·\underline{y} in it are neighbours. In “geometry based on the ring of dual numbers ℝ[ϵ]” (as in Hjelmslev’s [5]), one may define a ∼-relation, based on elements d of square 0, but the resulting monads will not be focused, essentially by the above argument. For, in [5], there is a fixed ϵ ∈ R such that every element d ∈ R with d^{2} = 0 is of the form ϵ · x.

We shall prove that in R^{n}, with metric derived from the standard inner product ⟨−, −⟩, the touching of spheres is focused. We first prove

8.5. Proposition. Let U ⊆ R^{n} be a finite dimensional linear subspace (meaning here that U is a linear direct summand in R^{n}); Then the following are equivalent for a vector \underline{a} ∈ R^{n} (assumed to be proper, i.e. \underline{a}#0)

1) \underline{a} ⊥ U

2) for all \underline{d} ∈ D(n) ∩ U, we have |\underline{a}| = |\underline{a} + \underline{d}|.

Proof. Since \underline{a} is proper, then so is \underline{a} + \underline{d}, so the two norms mentioned are > 0, so their equality is equivalent to the equality of their squares, i.e. to

⟨\underline{a}, \underline{a}⟩ = ⟨\underline{a} + \underline{d}, \underline{a} + \underline{d}⟩ = ⟨\underline{a}, \underline{a}⟩ + 2⟨\underline{a}, \underline{d}⟩.

So the Proposition says that \underline{a} ⊥ U iff ⟨\underline{a}, \underline{d}⟩ = 0 for all \underline{d} ∈ D(n) ∩ U; or, \underline{a} ⊥ U iff the linear functional ⟨\underline{a}, −⟩ : V → R restricts to 0 on D(n) ∩ U. The latter assertion is equivalent (like in the proof of Proposition 8.2) to saying that the functional restricts to 0 on all of U, i.e. to \underline{a} ⊥ U.

By suitable parallel translation, this implies, for U any finite dimensional affine subspace of R^{n}, that we have

8.6. Proposition. Let \underline{a} ∈ R^{n}. If \underline{b} is the foot (orthogonal projection) of \underline{a} on U, then \underline{a} has the same distance to all points \underline{b}′ ∈ U with \underline{b}′ ∼ \underline{b}; and conversely.

Now consider the special case where the affine subspace U is a hyperplane H, so of dimension n − 1. We consider in the following spheres A, whose center \underline{a} are apart from H, so dist(\underline{a}, \underline{x}) is defined for every \underline{x} ∈ H. Spheres A in R^{n} have dimension n − 1. Then one has (like in the Proposition 8.3) that if \underline{b} ∈ A∩H has 𝔐(\underline{b})∩H ⊆ A or 𝔐(\underline{b})∩A ⊆ H, then we have equality 𝔐(\underline{b})∩H = 𝔐(\underline{b})∩A, i.e. H and A touch each other in \underline{b}. Therefore

8.7. Proposition. 1) Let H be a hyperplane, and let a point \underline{a} (apart from H) have foot \underline{b} on H. Then H touches the sphere A := S(\underline{a}, r) in \underline{b}, where r = \underline{a}\underline{b}; 2) Conversely, if a sphere with center \underline{a} touches H in a point \underline{b}, then \underline{b} is the foot of \underline{a} on H.

Proof. For 1): Since \underline{b} is the foot of \underline{a}, we have \underline{a}\underline{b}′ = r for all \underline{b}′ ∈ H with \underline{b}′ ∼ \underline{b}, which is to say 𝔐(\underline{b}) ∩ H ⊆ S(\underline{a}, r), which as argued is the touching condition. – For 2), the assumption gives that 𝔐(\underline{b}) ∩ H ⊆ A, so all points \underline{b}′ ∈ H with \underline{b}′ ∼ \underline{b} have same distance to \underline{a}, so \underline{b} is the foot of \underline{a} on H, by Proposition 8.6 (the “conversely”-part).

8.8. Proposition. If A and B are spheres with centres \underline{a} and \underline{b}, respectively, then if \underline{x} and \underline{y} are in A ∩ B, ⟨\underline{x} − \underline{y}, \underline{a} − \underline{b}⟩ = 0

Proof. Let r and s be the radii of the two spheres. Then we have ⟨(\underline{x} − \underline{a}), (\underline{x} − \underline{a})⟩ = r^{2} and ⟨(\underline{x} − \underline{b}), (\underline{x} − \underline{b})⟩ = s^{2}, and similarly for \underline{y}. Then ⟨\underline{x} − \underline{y}, \underline{a} − \underline{b}⟩ = 0 follows by simple arithmetic.

Since feet (orthogonal projections) are unique, it follows that a sphere can touch a hyperplane in at most one point; and furthermore, the touching set is focused, being of the form 𝔐(b) ∩ H for a hyperplane in R^{n}, so is of the form D(n − 1). This proves the first assertion in

8.9. Corollary. In R^{n}, the touching of spheres with hyperplanes is focused. Also, the touching of two spheres (non-concentric in the sense that their centers are apart) is focused.

Proof. To prove the second assertion, let the spheres be A and C, with centers \underline{a} and \underline{c}, respectively (with \underline{a}#\underline{c}), and assume that A and C touch in a point \underline{b}. Let H be the hyperplane through \underline{b} and orthogonal to the line connecting \underline{a} and \underline{c}. Then \underline{b} is the foot of \underline{a} on H. So by Proposition 8.7, A touches H in \underline{b}. Similarly C touches H in \underline{b}, so

𝔐(\underline{b}) ∩ A = 𝔐(\underline{b}) ∩ H = 𝔐(\underline{b}) ∩ C,

and the middle set is known to be focused, since H is a hyperplane i R^{n}. We finally have to argue that the touching point \underline{b} is unique. If \underline{b}_{1} were another point in which the spheres touch, the hyperplane H_{1} through \underline{b}_{1} and orthogonal to the line from \underline{a} to \underline{c} is the same as H by Proposition 8.8, so \underline{b}_{1}, being the foot of \underline{a} on H_{1} = H, is the same as \underline{b}, proving the uniqueness of a possible touching point \underline{b}.

8.10. Remark. If A and H are as in Proposition 8.7, we have of course 𝔐(\underline{b}) ∩ H ⊆ A ∩ H, but we cannot conclude that 𝔐(\underline{b}) ∩ H equals A ∩ H; to wit, the unit sphere A in R^{3} with center (0, 0, 1) touches the xy-plane H with touching set D(2) × {0}, but A ∩ H = {(x, y, 0) | x^{2} + y^{2} = 0}, which is in general larger than D(2) × {0}.

The set {(x, y) ∈ R^{2} | x^{2} + y^{2} = 0} has been a puzzling “red herring” since the early days of SDG, see e.g. the Section “Loose Ends” in [6] ; in what sense is it an infinitesimal object? So the touching of spheres A and hyperplanes H in \underline{b} need not be “clean” in the sense that the touching set 𝔐(\underline{b}) ∩ H equals A ∩ H.

On the other hand, the unit circle A in R^{2} with center (0, 1) has A ∩ H = {(x, 0) | x^{2} = 0} = D × {0}; see the picture I.1.2 in [6]. So in R^{2}, “cleanness” can be asserted; in the preliminary version [10] it is incorrectly stated for general R^{n}.

References

[1] V.I. Arnold, Mathematical Methods of Classical Mechanics, Graduate Texts in Math. 60, Springer 1978.

[2] H. Busemann, On Spaces in Which Two Points Determine a Geodesic, Trans. Amer. Math. Soc. 54 (1943), 171-184.

[3] H. Busemann, Synthetische Differentialgeometrie, Jahresbericht DMV 71 (1969), 1- [4] H. Busemann, Recent Synthetic Differential Geometry, Ergebnisse der Mathematik

54, Springer 1970.

[5] J. Hjelmslev, Die natürliche Geometrie, Abhandlungen aus dem Math. Sem. der Uni. Hamburg (2), (1923), 1-36.

[6] A. Kock, Synthetic Differential Geometry, London Math. Soc. Lecture Notes 51, Cambridge U.P. 1981; 2nd Edition, London Math. Soc. Lecture Notes 333, Cambridge U.P. 2006.

[7] A. Kock, Envelopes – Notion and Definiteness, Beiträge zur Alg. und Geom. 48 (2007), 345-350.

[8] A. Kock, Synthetic Geometry of Manifolds, Cambridge Tracts in Math. 180, Cambridge U.P 2010.

[9] A. Kock, New Methods for Old Spaces: Synthetic Differential Geometry, arXiv 1610.00286. To appear in New Spaces in Mathematics and Physics, ed. M. Anel and G. Catren.

[10] A. Kock, Metric spaces and SDG (preliminary version), arXiv:1610.10005v3.

[11] R. Lavendhomme, Basic Concepts of Synthetic Differential Geometry, Kluwer 1996.

[12] S. Lie, Zur Theorie partieller Differentialgleichungen, Götting. Nachrichten 1872.

[13] S. Lie, Geometrie der Berührungstransformationen, Leipzig 1896 (reprint Chelsea Publ. Co., New York 1977).

[14] I. Moerdijk and G.E. Reyes, Models for Smooth Infinitesimal Analysis, Springer 1991.

Department of Mathematics, University of Aarhus, DK 8000 Aarhus C, Denmark

Email: kock@math.au.dk
This article may be accessed at http://www.tac.mta.ca/tac/

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available from the journal’s server at http://www.tac.mta.ca/tac/. It is archived electronically and in printed paper format.

Subscription information Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors The typesetting language of the journal is TEX, and L^{A}TEX2e is required. Articles in PDF format may be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors.

Clemens Berger, Université de Nice-Sophia Antipolis: cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown(at)btinternet.com Valeria de Paiva: Nuance Communications Inc: valeria.depaiva@gmail.com Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Kathryn Hess, Ecole Polytechnique Fédérale de Lausanne: kathryn.hess@epfl.ch Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, Macquarie University: steve.lack@mq.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@buffalo.edu Tom Leinster, University of Edinburgh: Tom.Leinster@ed.ac.uk Ieke Moerdijk, Utrecht University: i.moerdijk@uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Giuseppe Rosolini, Università di Genova: rosolini@disi.unige.it Alex Simpson, University of Ljubljana: Alex.Simpson@fmf.uni-lj.si James Stasheff, University of North Carolina: jds@math.upenn.edu Ross Street, Macquarie University: ross.street@mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Université du Québec à Montréal : tierney.myles4@gmail.com R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca